# Harmony in 10-TET: From Parallel Universes to the Desargues Configuration

### v1

@@ Abstract.
The 10-tone equal temperament (10-TET) constitutes a distinct harmonic environment with step sizes exceeding the standard semitone, yet it has historically remained on the fringes of music theory. This paper proposes a radical shift in perspective, treating $\mathbb{Z}_{10}$ not as an imperfect imitation of 12-TET, but as an independent harmonic universe with its own rigorous logic. We classify all possible harmonic systems in 10-TET using a single structural parameter $\Delta$, representing the difference between the harmonic mediants (the major and minor “thirds”) which sum to the generator (the “fifth”).
Our analysis reveals a rich landscape of possibilities. The system defined by $\Delta=1$, which is naturally the first to consider, initially appears to suffer from a fundamental structural flaw: its harmonic space fractures into two disjoint universes. Conversely, the “Wide” system ($\Delta=5$) emerges as a geometrically superior structure, isomorphic to the Desargues Configuration $(10_{3})_{1}$ and forming a fully connected, vertex-transitive graph. Between these extremes lies the “Tritone” system ($\Delta=3$), which forms a connected Decagonal Prism. Our main result is the discovery of a topological isomorphism between the “restored” Acoustic system ($\Delta=1$) and the Tritone system ($\Delta=3$): despite disjoint intervallic definitions, both generate the Generalized Petersen graph $GP(10,1)$. These results suggest that in 10-TET, the path to harmonic symmetry lies in the geometry of the projective plane and prism graphs rather than pure acoustics.
@@ 1. Introduction
The 10-tone equal temperament (10-TET) has long remained a theoretical curiosity, overshadowed by the ubiquity of the standard 12-tone system. While the system’s larger step sizes initially deterred classical adaptation, its theoretical foundations were established by Sethares [17], who demonstrated its unique capacity to support stable neutral intervals and novel timbral consonance. Practical interest in the system has steadily grown, with several compositions written for 10-TET, including some specifically composed for a 10-TET piano by Hunt [8], Senpai [16], Sevish [18], and Hideya [7]. However, while these works explored the electronic domain or software simulations, purely acoustic implementations remained elusive until the recent construction of the first decaphonic piano by renowned jazz pianist Możdżer and his collaborators [13] (see our Appendix A for more details). Despite these advances, 10-TET still lacks a comprehensive harmonic theory comparable to the rich structural tradition of Western tonal music. The goal of this paper is to provide a fully self-contained theory of possible harmonies within the system $\mathbb{Z}_{10}$. We aim to define precise analogues of Major and Minor chords, not through arbitrary selection, but via a rigorous generative process, and to construct the resulting harmonic networks—or Tonnetze—associated with them.
@@
This algorithmic approach allows us to scan through all possible values of $\Delta$ for a given $n$, exhaustively classifying every possible harmonic universe the temperament can support. While we focus here on $n=10$, this methodology provides a general blueprint for exploring harmonic topologies in any equal temperament. Crucially, this framework satisfies the correspondence principle vital to scientific theory generalization: a new broader theory must contain the established theory as a special case. Indeed, when we apply our $n$-TET harmonic proposal to the standard 12-TET ($n=12$), we recover the standard Western harmonic system, including the classical Euler-Riemann Tonnetz as one of the fundamental models.
@@
    The “Narrow” choice ($\Delta=1$), characterized by $(t,s)=(4,3)$. This corresponds to the most acoustically consonant intervals (closest to the frequency ratios 3:2 and 5:4).
@@
    The “Tritone” choice ($\Delta=3$), characterized by $(t,s)=(5,2)$. This serves as a “middle ground” solution utilizing the tritone.
@@
    The “Wide” choice ($\Delta=5$), characterized by $(t,s)=(6,1)$. This system utilizes larger, more abstract interval steps.
A surprising and central finding of this work is the topological divergence—and subsequent unification—of these systems. We demonstrate that the acoustically motivated “Narrow” system $(4,3)$ suffers from a fundamental degeneracy: it splits into two disjoint, parallel universes (the “Blue” and “Red” sets). Full modulation is possible only via non-standard “wormhole” transitions.
In contrast, the “Wide” system $(6,1)$, which we analyze in the latter part of this paper, emerges as a geometrically perfect structure. We prove that this system generates a highly connected, vertex-transitive graph isomorphic to the famous Desargues Configuration from projective geometry.
Finally, in Sections 9 and 10, we analyze the intermediate “Tritone” system ($\Delta=3$). We find that it forms a robust Decagonal Prism topology. Most notably, we prove a formal isomorphism between this Tritone system and the “Restored” Acoustic system. This reveals that the disjoint acoustic universes, when properly connected, share the exact same topological backbone as the tritone system: the Generalized Petersen Graph $GP(10,1)$. Thus, 10-TET offers two distinct paths to symmetry: the projective geometry of the Desargues configuration ($\Delta=5$) and the prismatic topology shared by the Acoustic and Tritone systems ($\Delta=1,3$).
@@ 5.3. Hypothesis 3: The “Neutral” Difference ($\Delta=3$)
The solution $(t=5,s=2)$ is mathematically valid. However, the "Major Third" $t=5$ corresponds to exactly half the octave ($\sqrt{2}$, or the tritone). A chord constructed this way would be $C=\{0,5,7\}$. While structurally interesting, a system based on the tritone as a consonance diverges radically from the acoustic principles of tertian harmony.
@@ 6.3. Musical Interpretation of Modal Degeneracy
6.4. The Degenerate Tonnetz as a Levi Graph
This identity fundamentally alters the topological construction of the Tonnetz. In 12-TET, the "Relative" transformation connects a chord to a distinct neighbor ($R:D_{r}\to M_{r-3}$), creating a graph where every vertex has degree 3. In 10-TET, the "Relative" transformation ($D_{r}\to M_{r+7}$) maps the chord to itself. Consequently, the edge corresponding to $R$ becomes a loop, and the functional degree of each vertex drops to 2.
To visualize this structure globally, we construct the **Levi graph** of the configuration. We treat the set of 10 major chords as "points" and the set of 10 minor chords as "lines". An incidence (edge) exists if a major chord transforms into a minor chord via a Parallel ($P$) or Leading-Tone ($L$) motion.
@@ 6.4. The Degenerate Tonnetz as a Levi Graph
Figure 6. The full Levi graph of the 10-TET Tonnetz arranged on a circle. The blue edges correspond to the connections between "even" chords, while the red edges connect "odd" chords. The interleaving of the nodes visually demonstrates the symmetry of the system.
@@ 7. The Functional Perspective: Restoring Connectivity in 3D
However, from a functional (Riemannian) perspective, a chord is defined not just by its pitch content, but by its root. In this view, the major triad $D_{0}=(0,4,7)$ rooted at 0 is distinct from the minor triad $M_{7}=(7,0,4)$ rooted at 7, despite comprising the same notes. The “Relative” transformation ($R$) is no longer an identity map, but a bijection – a wormhole – mapping a chord from the “even universe” to the “odd universe” (since $r$ and $r+7$ have opposite parity).
@@ 7.1. Musical Implications: The Phantom Modulation
7.2. The Combinatorial Structure: The $(10_{3})$ Configuration
By explicitly distinguishing between $D_{r}$ and $M_{r+7}$ and treating the identity relation as a valid edge ($R$), the system reveals a highly regular internal geometry. The interactions of $P$, $L$, and $R$ generate a unified graph that can be analyzed using the language of combinatorial design.
@@ 7.2.1. The Geometric Structure
This structure is a realization of a combinatorial $(10_{3})$ configuration. We can interpret the 10 Minor chords as “Points” and the 10 Major chords as “Lines”. The incidence relation is defined by our harmonic connections: each Line (Major chord) passes through exactly 3 Points (the Minors it connects to), and each Point (Minor chord) lies on exactly 3 Lines.
@@ 7.2.2. Visualization: The Levi Graphs
In Figure 10, we present the same graph in the standard sequential notation, where chords are arranged $D_{0},M_{0},D_{1},M_{1}\dots$ along a single circle. This layout highlights the symmetry of the configuration and clearly shows the alternating nature of the Blue and Red subnetworks.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$
Figure 10. The Connected Levi Graph in standard circular notation. Blue and Red nodes/edges denote the two universes; dashed gold lines denote the Identity Wormholes.
@@ 7.2.3. Minimal Chiral Cycles
The introduction of wormholes significantly reduces the girth (shortest cycle length) of the graph. While the intra-universe graphs have a girth of 10, the restored $(10_{3})$ graph allows for tight loops of length 4.
@@
In the restored $(4,3)$ system, there are exactly two isomorphism classes of minimal cycles (of length 4) under the rotational symmetry of $\mathbb{Z}_{10}$. These classes are distinguished by their chirality (direction of winding through the wormhole).
@@
These two sequences are non-isomorphic because they traverse the graph with opposite chirality, effectively winding around the toroidal structure of the configuration in opposite directions. ∎
@@
Figure 11. Visualisation of the two chiral minimal cycles. Straight edges connect the nodes, with arrows placed at the midpoint of each segment to indicate direction. These cycles represent the fundamental loops of the $(10_{3})$ configuration.
@@ 7.3. Algebraic Characterization and Graph Invariants
To properly classify the functional 10-TET graph within the family of cubic graphs, we must first establish its fundamental algebraic invariants. By inspection of the graph constructed in Section 7.2, we identify the following properties:

### v2

# Harmony in 10-TET: From Parallel Universes to Cousins of Desargues Configuration

@@ Abstract.
The 10-tone equal temperament (10-TET) constitutes a distinct harmonic environment with step sizes exceeding the standard semitone, yet it has historically remained on the fringes of music theory. This paper proposes a radical shift in perspective, treating 10-TET not as an imperfect imitation of 12-TET, but as an independent harmonic universe with its own rigorous logic. We classify all possible harmonic systems in 10-TET using a single structural parameter $\Delta$, representing the difference between the harmonic mediants (the major and minor “thirds”) which sum to the generator (the “fifth”).
Our analysis reveals a rich landscape of possibilities. The system defined by $\Delta=1$ initially appears to suffer from a fundamental structural flaw: its harmonic space fractures into two disjoint, parallel universes. Conversely, the “Wide” system ($\Delta=5$) emerges as a geometrically superior structure, isomorphic to the Unique Cyclic Configuration in the classification of combinatorial $10_{3}$ configurations, forming a fully connected, vertex-transitive graph. Between these extremes lies the “Tritone” system ($\Delta=3$), which topologically is similar to the $\Delta=1$ system, but it has a significantly larger symmetry group, due to the “Tritone” modal transformation. We conclude by providing a complete classification of these systems as bipartite circulant graphs, demonstrating that 10-TET is a mathematically complex and musically viable alternative to the standard 12-tone system.
@@ 1. Introduction
The 10-tone equal temperament (10-TET) has long remained a theoretical curiosity, overshadowed by the ubiquity of the standard 12-tone system. While the system’s larger step sizes initially deterred classical adaptation, its theoretical foundations were established by Sethares [19], who demonstrated its unique capacity to support stable neutral intervals and novel timbral consonance. Practical interest in the system has steadily grown, with several compositions written for 10-TET, including some specifically composed for a 10-TET piano by Hunt [12], Senpai [18], Sevish [20], and Hideya [11]. However, while these works explored the electronic domain or software simulations, purely acoustic implementations remained elusive until the recent construction of the first decaphonic piano by renowned jazz pianist Możdżer and his collaborators [16] (see our Appendix A for more details).
Despite these advances, 10-TET still lacks a comprehensive harmonic theory comparable to the rich structural tradition of Western tonal music. The goal of this paper is to provide a fully self-contained theory of possible harmonies within the system $\mathbb{Z}_{10}$. We aim to define precise analogues of Major and Minor chords, not through arbitrary selection, but via a rigorous generative process, and to construct the resulting harmonic networks—or Tonnetze—associated with them.
@@
This algorithmic approach allows us to scan through all possible values of $\Delta$ for a given $n$, exhaustively classifying every possible harmonic universe the temperament can support. While we focus here on $n=10$, this methodology provides a general blueprint for exploring harmonic topologies in any equal temperament.
Crucially, this framework satisfies the correspondence principle vital to scientific theory generalization: a new broader theory must contain the established theory as a special case. Indeed, when we apply our $n$-TET harmonic proposal to the standard 12-TET ($n=12$), we recover the standard Western harmonic system, including the classical Euler-Riemann Tonnetz as one of the fundamental models.
@@
    The “Narrow”(or “Acoustic”) choice ($\Delta=1$), characterized by $(t,s)=(4,3)$. This corresponds to the first–to–try intervals but suffers from modal degeneracy and topological splitting into two disjoint universes.
@@
    The “Tritone” choice ($\Delta=3$), characterized by $(t,s)=(5,2)$, which serves as a highly symmetric “middle ground” using tritone.
@@
    The “Wide” choice ($\Delta=5$), characterized by $(t,s)=(6,1)$, which generates a single, connected component isomorphic to a one of the cyclic configurations in combinatorial geometry. This system utilizes the largest interval steps.
A surprising finding of this work is the topological divergence—and subsequent unification—of the acoustically motivated “Narrow” system $(4,3)$. It suffers from a fundamental degeneracy, spliting into two disjoint, parallel universes (the “Blue” and “Red” sets). Full modulation in this system is possible only via non-standard “wormhole” transitions.
In contrast, the “Wide” system $(6,1)$, which we analyze in the latter part of this paper, emerges as a geometrically perfect structure. We prove that this system generates a highly connected, vertex-transitive graph isomorphic to the unique cyclic configuration $(10_{3})_{3}$ in census [3]. This structure stands apart from the famous Desargues $10_{3}$ configuration (which is not cyclic) as a singular intersection between the axioms of combinatorial geometry and the cyclic symmetry inherent in musical theory.
By exploring these systems, we demonstrate that 10-TET is not merely a microtonal experiment, but a structurally rich environment where the choice of intervals fundamentally dictates the topology of musical space. The paper is organized as follows: Section 2 motivates our choice of the ‘fifth’ $q$, which is crucial for obtaining possible candidates fo 10-TET harmonies; Section 3 formalizes the definition of chords; and Sections 4 through 10 provide a detailed graph-theoretical analysis of each harmonic candidate and their respective symmetry groups.
@@ 5.3. Hypothesis 3: The “Neutral” Difference ($\Delta=3$)
The solution $(t=5,s=2)$ is mathematically valid. However, the "Major Third" $t=5$ corresponds to exactly half the octave ($\sqrt{2}$, or the tritone). A chord constructed this way would be $C=\{0,5,7\}$.
@@ 6.3. Musical Interpretation of Modal Degeneracy
6.4. The Degenerate Tonnetz and its Levi Graph
To understand the topology of this system, we introduce the standard Neo-Riemannian transformations adapted to 10-TET. For every Major chord $D_{r}$ and Minor chord $M_{r}$, we define the operations $P$ (Parallel), $L$ (Leading-Tone), and $R$ (Relative) as follows:
  • Parallel ($P$): Preserves the root and fifth.
    $P(D_{r})=M_{r},\quad P(M_{r})=D_{r}$
  • Leading-Tone ($L$): Preserves the minor third (Major chord) or major third (Minor chord).
    $L(D_{r})=M_{r+4},\quad L(M_{r})=D_{r-4}$
  • Relative ($R$): Preserves the major third (Major chord) or minor third (Minor chord).
    $R(D_{r})=M_{r-3},\quad R(M_{r})=D_{r+3}$
All indices are taken modulo 10. These transformations map the chords as shown in the tables below. Note that the $R$ transformation is crossed out because, in this specific system, it acts as the identity operation on the set of chords.
Input $P$ $L$ $R$
$D_{0}$ $M_{0}$ $M_{4}$ $M_{7}$
$D_{1}$ $M_{1}$ $M_{5}$ $M_{8}$
$D_{2}$ $M_{2}$ $M_{6}$ $M_{9}$
$D_{3}$ $M_{3}$ $M_{7}$ $M_{0}$
$D_{4}$ $M_{4}$ $M_{8}$ $M_{1}$
$D_{5}$ $M_{5}$ $M_{9}$ $M_{2}$
$D_{6}$ $M_{6}$ $M_{0}$ $M_{3}$
$D_{7}$ $M_{7}$ $M_{1}$ $M_{4}$
$D_{8}$ $M_{8}$ $M_{2}$ $M_{5}$
$D_{9}$ $M_{9}$ $M_{3}$ $M_{6}$
Figure 6. Transformations on Major Chords. The $R$ column is crossed out because $M_{r-3}\equiv D_{r}$ is a trivial identity.
Input $P$ $L$ $R$
$M_{0}$ $D_{0}$ $D_{6}$ $D_{3}$
$M_{1}$ $D_{1}$ $D_{7}$ $D_{4}$
$M_{2}$ $D_{2}$ $D_{8}$ $D_{5}$
$M_{3}$ $D_{3}$ $D_{9}$ $D_{6}$
$M_{4}$ $D_{4}$ $D_{0}$ $D_{7}$
$M_{5}$ $D_{5}$ $D_{1}$ $D_{8}$
$M_{6}$ $D_{6}$ $D_{2}$ $D_{9}$
$M_{7}$ $D_{7}$ $D_{3}$ $D_{0}$
$M_{8}$ $D_{8}$ $D_{4}$ $D_{1}$
$M_{9}$ $D_{9}$ $D_{5}$ $D_{2}$
Figure 7. Transformations on Minor Chords. The $R$ column is crossed out because $D_{r+3}\equiv M_{r}$ is a trivial identity.
This identity fundamentally alters the topological construction of the Tonnetz. In 12-TET, the “Relative” transformation connects a chord to a distinct neighbor ($R:D_{r}\to M_{r-3}$), creating a graph where every vertex has degree 3. In 10-TET, the “Relative” transformation ($D_{r}\to M_{r+7}$) maps the chord to itself. Consequently, the edge corresponding to $R$ becomes a loop, and the functional degree of each vertex drops to 2.
To visualize this structure globally, we construct the **Levi graph** of the configuration. We treat the set of 10 major chords as “points” and the set of 10 minor chords as “lines”. An incidence (edge) exists if a major chord transforms into a minor chord via a Parallel ($P$) or Leading-Tone ($L$) motion.
@@ 6.4. The Degenerate Tonnetz and its Levi Graph
Figure 8. The full Levi graph of the 10-TET Tonnetz arranged on a circle. The blue edges correspond to the connections between “even” chords, while the red edges connect “odd” chords. The interleaving of the nodes visually demonstrates the symmetry of the system.
@@ 7. The Functional Perspective: Restoring Connectivity in 3D
However, from a functional (Riemannian) perspective, a chord is defined not just by its pitch content, but by its root. In this view, the major triad $D_{0}=\{0,4,7\}$ rooted at 0 is distinct from the minor triad $M_{7}=\{7,0,4\}$ rooted at 7, despite comprising the same notes. The “Relative” transformation ($R$) is no longer an identity map, but a bijection – a wormhole – mapping a chord from the “even universe” to the “odd universe” (since $r$ and $r+7$ have opposite parity).
@@
We explicitly list these restored functional connections in the tables below. Unlike the degenerate case, the $R$-column now provides valid, distinct outputs that link the previously disjoint cycles.
Input $P$ $L$ $R$
$D_{0}$ $M_{0}$ $M_{4}$ $M_{7}$
$D_{1}$ $M_{1}$ $M_{5}$ $M_{8}$
$D_{2}$ $M_{2}$ $M_{6}$ $M_{9}$
$D_{3}$ $M_{3}$ $M_{7}$ $M_{0}$
$D_{4}$ $M_{4}$ $M_{8}$ $M_{1}$
$D_{5}$ $M_{5}$ $M_{9}$ $M_{2}$
$D_{6}$ $M_{6}$ $M_{0}$ $M_{3}$
$D_{7}$ $M_{7}$ $M_{1}$ $M_{4}$
$D_{8}$ $M_{8}$ $M_{2}$ $M_{5}$
$D_{9}$ $M_{9}$ $M_{3}$ $M_{6}$
Figure 10. Functional Transformations on Major Chords. The $R$ transformation is now active, mapping $D_{r}$ to the functionally distinct $M_{r-3}$.
Input $P$ $L$ $R$
$M_{0}$ $D_{0}$ $D_{6}$ $D_{3}$
$M_{1}$ $D_{1}$ $D_{7}$ $D_{4}$
$M_{2}$ $D_{2}$ $D_{8}$ $D_{5}$
$M_{3}$ $D_{3}$ $D_{9}$ $D_{6}$
$M_{4}$ $D_{4}$ $D_{0}$ $D_{7}$
$M_{5}$ $D_{5}$ $D_{1}$ $D_{8}$
$M_{6}$ $D_{6}$ $D_{2}$ $D_{9}$
$M_{7}$ $D_{7}$ $D_{3}$ $D_{0}$
$M_{8}$ $D_{8}$ $D_{4}$ $D_{1}$
$M_{9}$ $D_{9}$ $D_{5}$ $D_{2}$
Figure 11. Functional Transformations on Minor Chords. The $R$ transformation is now active, mapping $M_{r}$ to the functionally distinct $D_{r+3}$.
@@ 7.1. Musical Implications: The Phantom Modulation
7.2. The Combinatorial Structure: The Cubic Graph
By explicitly distinguishing between $D_{r}$ and $M_{r+7}$ and treating the identity relation as a valid edge ($R$), the system reveals a highly regular internal geometry. The interactions of $P$, $L$, and $R$ generate a unified graph that can be analyzed using the language of graph theory.
@@ 7.2.1. The Geometric Structure
This structure constitutes a connected cubic bipartite graph. We can interpret the 10 Minor chords as “Points” and the 10 Major chords as “Lines”. The incidence relation is defined by our harmonic connections: each Line (Major chord) passes through exactly 3 Points (the Minors it connects to), and each Point (Minor chord) lies on exactly 3 Lines.
However, it should be mentioned that these are not geometric points and lines in the strict sense. As we will see in the Section 7.2.3, in this interpretation the graph admits that two points lie on two different lines and two lines can meet in more than one point.
@@ 7.2.2. Visualization: The Levi Graphs
In Figure 14, we present two visualizations of the graph. On the left, chords are arranged in the standard sequential notation ($D_{0},M_{0},D_{1},M_{1}\dots$), highlighting the local alternating nature of the Blue and Red subnetworks. On the right, the chords are arranged according to the Hamiltonian cycle (see Section 7.3) found in the system ($D_{0},M_{0},D_{3},M_{3}\dots$), revealing the global symmetry of the connections.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$(a) Standard Sequence
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$(b) Hamiltonian Sequence
Figure 14. The Connected Levi Graph visualized in two arrangements. (a) Standard sequential notation. (b) Hamiltonian perimeter arrangement. Blue and Red nodes denote the parity of the root; dashed gold lines denote the Identity Wormholes connecting them.
@@ 7.2.3. Minimal Chiral Cycles
The introduction of wormholes significantly reduces the girth (shortest cycle length) of the graph. While the intra-universe graphs have a girth of 10, the restored cubic graph allows for tight loops of length 4.
@@
In the restored cubic system, there are exactly two isomorphism classes of minimal cycles (of length 4) under the rotational symmetry of $\mathbb{Z}_{10}$. These classes are distinguished by their chirality (direction of winding through the wormhole).
@@
These two sequences are non-isomorphic because they traverse the graph with opposite chirality, effectively winding around the toroidal structure of the graph in opposite directions. ∎
@@
Figure 15. Visualisation of the two chiral minimal cycles. Straight edges connect the nodes, with arrows placed at the midpoint of each segment to indicate direction. These cycles represent the fundamental loops of the cubic graph.
@@ 7.3. Algebraic Characterization and Graph Invariants
To properly classify the functional 10-TET graph within the family of cubic graphs, we must first establish its fundamental algebraic invariants. By inspection of the graph constructed in Section 7.2, we identify the following properties (more of them will be listed in Section LABEL:sec:algebraic_invariants):
@@
    Remark: The existence of these 4-cycles implies that in the point-line interpretation, two distinct "lines" (e.g., $D_{0}$ and $D_{3}$) intersect at two distinct "points" (e.g., $M_{0}$ and $M_{7}$). Consequently, two distinct points can determine more than one line. This violates the axioms of a linear configuration, distinguishing this graph from the rigid geometric structures discussed later.